# Measurement of Sudden Beam Loss Events Using Bunch-by-Bunch BPMs at SuperKEKB

### v1

@@ IV Analysis Using Multiple Monitoring Systems
IV.1 Relationship between pressure burst locations and bunch oscillations
We begin by describing the phenomenon of pressure bursts [6], which are suspected to be closely related to SBL events. During SBL events, a sudden and abnormal increase in vacuum pressure inside the beam chamber is frequently observed at certain points in the main ring. As an example, Fig. 12 shows the time evolution of the vacuum pressure. It was measured by the CCG installed at D10_L02 during the SBL event shown in Fig. 2. It can be seen that the vacuum pressure rises sharply after the beam current drops to zero due to the beam abort. Because the response time of the CCG is relatively slow, it is believed that the pressure spike occurs either simultaneously with or just before the beam abort. We refer to such phenomena as “pressure bursts.” In general, pressure bursts are rarely observed during controlled, safe beam aborts that do not involve beam loss.
@@ IV.1 Relationship between pressure burst locations and bunch oscillations
The duration of bunch oscillations prior to the beam abort is determined using the method described in Section III.3. The results are categorized by pressure burst location, and the corresponding histograms are shown in Fig. 13. Focusing on the two most frequently observed pressure burst locations—D10_L02/03 and D10_L05—we find the following:
@@
IV.2 Potential scenario of SBL evolution arising at D10_L02/L03
@@
If the beam receives a kick in the D10 section, the resulting oscillation would propagate downstream, first passing the Fuji-RFSoC and then traveling through Section 1 to the D5-RFSoC. Naively, this would result in simultaneous observation of oscillation onset at both BORs in these plots. However, the observed amplitude at each location depends on the betatron phase relative to the kick location. This may delay the detection of the oscillation onset under conditions where position oscillation is not visible. As inferred earlier, the beam likely received the kick force near the midpoint between D10_L02 and D10_L03. The bunch position observed at downstream “location 2” after receiving a kick at “location 1” can be expressed as:
$y_{2}\approx\sqrt{\beta_{y1}\beta_{y2}}\sin\Psi_{12}\Delta y_{1}^{\prime}$ (2)
Here, $y_{2}$ is the bunch position at location 2, $\beta_{y1},\beta_{y2}$ are the beta functions at locations 1 and 2, $\Psi_{12}$ is the betatron phase advance, and $\Delta y_{1}^{\prime}$ is the kick angle. From this expression, the oscillation is most visible when $\Psi_{12}\approx(n+0.5)\pi$, and minimized when $\Psi_{12}\approx n\pi$, where $n$ is an integer. Table 6 summarizes the betatron phases of the Fuji-RFSoC and the D5-RFSoC relative to the midpoint between D10_L02 and D10_L03, based on Table 3. We find that the Fuji-RFSoC is approximately at a half-integer multiple of $\pi$, while the D5-RFSoC is near an integer multiple for both horizontal and vertical directions. Thus, if a kick occurs at the midpoint, it is more likely to be strongly visible at the Fuji-RFSoC, consistent with the earlier onset detection there.
@@
Horizontal betatron phase [rad/$\pi$] 0 21.40 40.89
Vertical betatron phase [rad/$\pi$] 0 22.56 42.93
@@
IV.3 Potential scenario of SBL evolution arising at D10_L05
@@
To understand the sequence of events, we analyze which collimator in the D06 section the bunch hits first, based on fast loss monitor signals. In addition to the loss monitors for beam abort described in Section II.3, several fast-response loss monitors dedicated to SBL observation are installed near collimators [14]. These include cesium iodide scintillator with photomultiplier tube (CsI+PMT) monitors at D02V1 and D06V2, and electron multiplier tube (EMT) monitors at D06H3, D06H4, D06V1, D05V1, and D03H1. Figure 19 shows the loss monitor signals recorded during the event in Fig. 17 [15, 14]. The blue dashed vertical lines indicate the same timing as the turn axis in Fig. 17. According to this figure, the loss monitor installed at D06V1 detected radiation earliest—approximately 2.5 turns before the beam abort trigger. This closely follows the onset of horizontal oscillation observed at the Fuji-RFSoC. It thus appears that the bunch, already oscillating at the Fuji-RFSoC, first collided with the vertical collimator D06V1. Although D06H3 and D06H4 collimators (horizontal collimators located upstream of D06V1) also have loss monitors; they detected losses one turn later. This pattern was consistently observed in all three additional D10_L05-related events. These results suggest that vertical oscillations were likely initiated simultaneously with horizontal oscillations, and it was the vertical oscillation that first caused beam loss at D06V1.
@@
Horizontal betatron phase [rad/$\pi$] 0 20.73 31.23
Vertical betatron phase [rad/$\pi$] 0 22.01 33.44
@@
Figure 20 shows the beta and dispersion functions near D10_L05. A vertical focusing quadrupole magnet is located at D10_L05, causing the vertical beta function to peak. According to Eq. (2), the larger the beta function at the kick location, the greater the resulting oscillation. Thus, a kick at D10_L05 would significantly enhance vertical oscillation, potentially leading to immediate impact with vertical collimators and charge loss.
@@
V Summary
This study has significantly advanced our understanding of SBL and has provided crucial insights for future SBL mitigation strategies and luminosity improvement efforts.
Specifically, observations using the BOR revealed that bunch charge loss during SBL events predominantly occurs at the collimators in the D06 section. This finding suggests that substantial beam position oscillations likely occur before the beam enters the D06 section. Moreover, classifying SBL events based on the location of associated pressure bursts uncovered distinct patterns in the onset of oscillations. Further analysis from the viewpoint of the accelerator lattice implies a causal relationship wherein the beam receives a disturbance at the pressure burst location, subsequently leading to the development of SBL.
Going forward, a key challenge will be to generalize the interpretation of SBL evolution by applying the scenario investigated at each pressure burst site to other locations around the ring. By increasing the number and resolution of BORs, it will be possible to determine whether SBL is a location-independent phenomenon governed by universal parameters such as the beta or dispersion function, or whether it is strongly influenced by site-specific conditions not captured by lattice optics alone. Additionally, we will also focus on the possible beam size growth, as suggested in Appendix B, to deepen our understanding of the mechanism of SBL events.
Acknowledgements.
@@
Appendix A Discussion on duration of bunch position oscillation
In Section III.3, we discussed the duration of bunch position oscillations during SBL events. Here, we further examine this topic by comparing the observed oscillation durations with the results of previous tests.
During the development of the BOR, we conducted performance tests using the main ring’s bunch-by-bunch feedback system. Details of this test are reported in Ref. [5]. In the test, we reversed the phase of the feedback kicker in the bunch-by-bunch feedback system [16] to deliberately amplify the bunch oscillations in a total of 393 bunches, and monitored the resulting motion using the BOR. The feedback kicker was installed in the Fuji straight section. The results of this test are shown in Fig. 21.
@@ Appendix A Discussion on duration of bunch position oscillation
This figure demonstrates that BOR successfully captured the increase in bunch oscillation amplitude caused by the inverted kicker phase. A beam abort occurred after the right edge of the plot due to the excessive oscillation amplitude. The optics used in this test was the $\beta_{y}^{\ast}=1~{}\mathrm{mm}$ configuration, which features a relatively tight physical aperture. The strength of the kick applied in the test was equivalent to that applied under normal operating conditions, since the feedback gain settings remained unchanged. Even with deliberate destabilization using the feedback kicker, it took several hundred microseconds to induce sufficient beam loss to trigger a beam abort.
In contrast, as shown in Section III.3, SBL events typically cause beam aborts after only tens to about 100 µs of bunch position oscillation. Although it had been hypothesized that the feedback system itself might inadvertently induce SBL due to inadequate optimization, the results above suggest this scenario is unlikely. To induce beam loss within less than 100 µs, a much stronger force than the standard feedback kicker must act on the beam.
Appendix B Discussion on amplitude of bunch position oscillation
We revisit the feedback test discussed in Appendix A to consider the amplitude of position oscillations observed. According to Fig. 21, the bunch oscillations grew to approximately $\pm$0.6 mm (i.e., oscillation amplitude $\sim$1.2 mm). During the 750 µs interval over which the oscillation amplitude increased, no beam loss occurred, and measurement was uninterrupted. This indicates that, even under tight aperture optics with $\beta_{y}^{\ast}=1~{}\mathrm{mm}$, oscillations up to roughly 1.2 mm are tolerable.
However, as shown in Fig. 11, oscillation amplitudes observed during SBL events are generally smaller than this. This implies that, despite the oscillation amplitude being within the physically allowed aperture, substantial charge loss and beam aborts frequently occur during SBL events. This discrepancy suggests that beam loss during SBL may not be due solely to increased bunch position oscillations. Rather, it may result from a combination of position oscillation and simultaneous beam size growth. Indeed, beam size monitors have recorded signs of increased beam size during several SBL events [17], supporting this interpretation.

### v2

@@ 4 Analysis Using Multiple Monitoring Systems
4.1 Relationship between pressure burst locations and bunch oscillations
We begin by describing the phenomenon of pressure bursts [20], which are suspected to be closely related to SBL events. During SBL events, a sudden and abnormal increase in vacuum pressure inside the beam chamber is frequently observed at certain points in the main ring. As an example, Fig. 12 shows the time evolution of the vacuum pressure. It was measured by the CCG installed at D10_L02 during the SBL event shown in Fig. 2. It can be seen that the vacuum pressure rises sharply after the beam current drops to zero due to the beam abort. Because the response time of the CCG is relatively slow, it is believed that the pressure spike occurs either simultaneously with or just before the beam abort. We refer to such phenomena as “pressure bursts.” Note that, in general, pressure bursts are not observed during controlled, safe beam aborts that do not involve beam loss.
@@ 4.1 Relationship between pressure burst locations and bunch oscillations
The duration of bunch oscillations prior to the beam abort is determined using the method described in Section 3.3. The results are categorized by pressure burst location, and the corresponding histograms are shown in Fig. 13. Focusing on the two most frequently observed pressure burst locations—D10_L02/03 and D10_L05—we find the following:
@@
4.2 Potential scenario of SBL evolution arising at D10_L02/L03
@@
If the beam receives a kick in the D10 section, the resulting oscillation would propagate downstream, first passing the Fuji-RFSoC and then traveling through Section 1 to the D5-RFSoC. One might expect this to result in simultaneous observation of oscillation onset at both BORs in these plots. However, the observed amplitude at each location depends on the betatron phase relative to the kick location. This may delay the detection of the oscillation onset under conditions where position oscillation is not visible. As inferred earlier, the beam likely received the kick force near the midpoint between D10_L02 and D10_L03. The bunch position observed at downstream “location 2” after receiving a kick at “location 1” can be expressed as:
$y_{2}\approx\sqrt{\beta_{y1}\beta_{y2}}\sin\Psi_{12}\Delta y_{1}^{\prime},$ (4.1)
where $y_{2}$ is the bunch position at location 2, $\beta_{y1},\beta_{y2}$ are the beta functions at locations 1 and 2, $\Psi_{12}$ is the betatron phase advance between location 1 and 2, and $\Delta y_{1}^{\prime}$ is the kick angle. From this expression, the oscillation is most visible when $\Psi_{12}\approx(n+0.5)\pi$, and minimized when $\Psi_{12}\approx n\pi$, where $n$ is an integer. Table 6 summarizes the betatron phases of the Fuji-RFSoC and the D5-RFSoC relative to the midpoint between D10_L02 and D10_L03, based on Table 2. We find that the Fuji-RFSoC is approximately at a half-integer multiple of $\pi$, while the D5-RFSoC is near an integer multiple for both horizontal and vertical directions. Thus, if a kick occurs at the midpoint, it is more likely to be strongly visible at the Fuji-RFSoC, consistent with the earlier onset detection there.
@@
[…]
@@
4.3 Potential scenario of SBL evolution arising at D10_L05
@@
To understand the sequence of events, we analyze which collimator in the D06 section the bunch hits first, based on fast loss monitor signals. In addition to the loss monitors for beam abort described in Section 2.3, several fast-response loss monitors dedicated to SBL observation are installed near collimators [23]. These include cesium iodide scintillator with photomultiplier tube (CsI+PMT) monitors at D02V1 and D06V2, and electron multiplier tube (EMT) monitors at D06H3, D06H4, D06V1, D05V1, and D03H1. Figure 19 shows the loss monitor signals recorded during the event in Fig. 17 [16, 23]. The blue dashed vertical lines indicate the same timing as the turn axis in Fig. 17. According to this figure, the loss monitor installed at D06V1 detected radiation earliest—approximately 2.5 turns before the beam abort trigger. This closely follows the onset of horizontal oscillation observed at the Fuji-RFSoC. It thus appears that the bunch, already oscillating at the Fuji-RFSoC, first collided with the vertical collimator D06V1. Although D06H3 and D06H4 collimators (horizontal collimators located upstream of D06V1) also have loss monitors; they detected losses one turn later. This pattern was consistently observed in all three additional D10_L05-related events. These results suggest that vertical oscillations were likely initiated simultaneously with horizontal oscillations, and it was the vertical oscillation that first caused beam loss at D06V1.
@@
[…]
@@
Figure 20 shows the beta and dispersion functions near D10_L05. A vertical focusing quadrupole magnet is located at D10_L05, causing the vertical beta function to peak. According to Eq. (4.1), the larger the beta function at the kick location, the greater the resulting oscillation. Thus, a kick at D10_L05 would significantly enhance vertical oscillation, potentially leading to immediate impact with vertical collimators and charge loss.
@@
5 Summary
This study has significantly advanced our understanding of SBL and provided crucial insights for future SBL mitigation strategies and luminosity improvement efforts.
Specifically, observations using the BOR revealed that bunch charge loss during SBL events predominantly occurs at the collimators in the D06 section. This finding suggests that substantial beam position oscillations likely occur before the beam enters the D06 section. Moreover, classifying SBL events based on the location of associated pressure bursts uncovered distinct patterns in the onset of oscillations. Further analysis from the viewpoint of the accelerator lattice implies a causal relationship wherein the beam receives a disturbance at the pressure burst location, subsequently leading to the development of SBL. These results clearly demonstrate the effectiveness of the BOR as a diagnostic instrument and establish a foundation for future studies of SBL phenomena in nanobeam accelerators.
Going forward, a key challenge will be to generalize the interpretation of SBL evolution by applying the scenario investigated at each pressure burst site to other locations around the ring. By increasing the number and resolution of BORs, it will be possible to determine whether SBL is a location-independent phenomenon governed by universal parameters such as the beta or dispersion function, or whether it is strongly influenced by site-specific conditions not captured by lattice optics alone. Additionally, we will also focus on the possible beam size growth, as suggested in Appendix A, to deepen our understanding of the mechanism of SBL events.
In addition, compared with similar fast beam-loss phenomena at the LHC, known as UFOs, SBL at SuperKEKB proceeds on an even shorter timescale. At the LHC, machine damage was mitigated by optimizing the beam loss monitor system [1]; however, for SBL such mitigation is insufficient, as a beam abort based solely on loss monitors cannot be executed quickly enough. This highlights the necessity of diagnostics with higher time resolution. In this work, by focusing on bunch-by-bunch beam position measurements, we were able to directly observe the onset of bunch oscillations. Furthermore, by combining these observations with pressure data, we found that the bunch oscillations are closely linked to pressure bursts. These findings represent entirely new insights and may provide valuable observables not only for SuperKEKB but also for LHC and future facilities such as proposed Higgs factories. Nevertheless, whether SBL is a dust-induced beam instability, analogous to UFOs at the LHC, cannot be determined from the present results alone. To address this question, we plan to combine bunch position, pressure, and future beam size measurements with detailed simulations in order to move closer to the underlying mechanism of SBL.
6 Acknowledgements
@@
Appendix A Discussion on the amplitude of bunch position oscillations
In Section 3.4, we discussed the amplitude of bunch position oscillations during SBL events. Here, we further investigate this topic by comparing the observed oscillation amplitudes with the results of a previous test.
During the development of the BOR, we conducted a performance test using the main ring’s bunch-by-bunch feedback system. Details of this test are reported in Ref. [13]. In this test, we reversed the phase of the feedback kicker in the bunch-by-bunch feedback system [22] to deliberately amplify the bunch oscillations in a total of 393 bunches, and monitored the resulting motion using the BOR. The feedback kicker was installed in the Fuji straight section. The results of this test are shown in Fig. 21. This figure demonstrates that BOR successfully captured the increase in bunch oscillation amplitude caused by the inverted kicker phase. A beam abort occurred after the right edge of the plot due to the excessive oscillation amplitude. The optics used in this test was the $\beta_{y}^{\ast}=1~\mathrm{mm}$ configuration, which features a relatively tight physical aperture. The strength of the kick applied in the test was equivalent to that applied under normal operating conditions, since the feedback gain settings remained unchanged.
@@ Appendix A Discussion on the amplitude of bunch position oscillations
According to Fig. 21, the bunch oscillations grew to approximately $\pm$0.6 mm (i.e., oscillation amplitude $\sim$1.2 mm). During the 750 µs interval over which the oscillation amplitude increased, no beam loss occurred, and measurement was uninterrupted. This indicates that, even under tight aperture optics with $\beta_{y}^{\ast}=1~\mathrm{mm}$, oscillations up to roughly 1.2 mm are tolerable. However, as shown in Fig. 11, oscillation amplitudes observed during SBL events are generally smaller than this. This implies that, despite the oscillation amplitude being within the physically allowed aperture, substantial charge loss and beam aborts frequently occur during SBL events. This discrepancy suggests that beam loss during SBL may not be due solely to increased bunch position oscillations. Rather, it may result from a combination of position oscillation and simultaneous beam size growth. Indeed, beam size monitors have recorded signs of increased beam size during several SBL events [11], supporting this interpretation.
In Fig. 21, a clear coherent growth of bunch position oscillations is observed, showing a pattern in which the oscillation appears to diffuse along the bunch train. In contrast, in several SBL events presented in this paper, the oscillation pattern along the bunch train is significantly different: the oscillations often appear as if the motion propagates along successive bunches. One possible explanation is that the feedback kicker applies kicks to each bunch in varying directions, whereas the kick source responsible for SBL may act over a certain period of time, continuously affecting multiple bunches. These observations suggest that the feedback system, including the feedback kicker, is unlikely to be the origin of SBL.